OpenStudy (anonymous):

Determine the slope of the line which passes through these two points: (-1, 3) and (6, 9).

OpenStudy (anonymous):

6/7

OpenStudy (anonymous):

y=ax+b a=(9-3)/(6-(-1))=6/7

OpenStudy (anonymous):

(9-3)/(6--1) 6/7

OpenStudy (anonymous):

Line equation = \[y-y _{0}=2a(x-x _{0})\]

OpenStudy (anonymous):

y−y 0 =a(x−x 0 ) this is right

OpenStudy (anonymous):

y-9=2*(6/7)*(x-6)

OpenStudy (anonymous):

don't mix it up with physic

OpenStudy (anonymous):

you're wrong amir.sat

OpenStudy (anonymous):

you're wrong

OpenStudy (anonymous):

not me

OpenStudy (anonymous):

review your lesson more timea lol

OpenStudy (anonymous):